E6-3 Barnes Company reports the following operating results for the month of August: sales Compute net income under different alternatives (units 5,000): variable costs $210,000; and fixed costs $75,000. M $325,000 independent courses of action to increase net income. anagement is considering the following Increase selling price by 10% with no change in total variable costs or sales volume. Reduce variable costs to 58% of sales 2 3 Instructions Compute the net income to be earned under each alternative. Which Reduce fixed costs by $15,000. course of action will produce the highest net income? mber in cells requesting a value; enter either a number or a formula in cells with a? Increase selling price by 10% with no change in total variable costs or sales volume. 2 Current selling price New selling price Round to nearest cent) Total sales Less: variable costs Contribution margin Less: fixed costs Net income Value Reduce variable costs to 58% of sales Total Sales Less: variable costs Contribution margin Less: fixed costs 2. Value Value ValueNet income Reduce fixed costs by $15,000 Total Sales Less: variable costs Contribution margin Less: fixed costs Net income 3 Value Value Value After you have completed E6-3, consider the following additional questions. ssume that unit selling price increased 5% with no change in total variable costs or sales volume. Assume variable costs decreased to 53% of sales Assume that fixed costs increased by $20,000. Which course of action will produce the highest net income?
Answer #1
Ans. Current situation :
Total Per unit
Sales 325000 65
Less: variable expenses 210000 42
Contribution margin 115000 23
Less: Fixed expenses 75000
Net income 40000
*Selling price per unit = Total sales / No. of units sold
325000 / 5000   =    $65 per unit
*Variable cost per unit = Total variable cost / No. of units sold
210000 / 5000    =   $42 per unit
E6 - 3
Ans. 1 New selling price =   65 * 110% 71.5
Total Per unit
Sales (5000*71.5) 357500 71.5
Less: variable expenses 210000 42
Contribution margin 147500 29.5
Less: Fixed expenses 75000
Net income 72500
Ans.2 New variable cost = 325000 * 58%   =     188500
Total Per unit
Sales 325000 65
Less: variable expenses 188500 37.7
Contribution margin 136500 27.3
Less: Fixed expenses 75000
Net income 61500
Ans.3 New fixed cost = (75000 - 15000) = 60000
Total Per unit
Sales 325000 65
Less: variable expenses 210000 42
Contribution margin 115000 23
Less: Fixed expenses 60000
Net income 55000
* The first course of action (Ans.1) will provide the highest net income.
*Additional questions:
Ans. 1 New selling price =   65 * 105% 68.25
Total Per unit
Sales (5000*71.5) 341250 68.25
Less: variable expenses 210000 42
Contribution margin 131250 26.25
Less: Fixed expenses 75000
Net income 56250
Ans.2 New variable cost = 325000 * 53%   =     172250
Total Per unit
Sales 325000 65
Less: variable expenses 172250 34.45
Contribution margin 152750 30.55
Less: Fixed expenses 75000
Net income 77750
Ans.3 New fixed cost = (75000 + 20000) = 95000
Total Per unit
Sales 325000 65
Less: variable expenses 210000 42
Contribution margin 115000 23
Less: Fixed expenses 95000
Net income 20000
* The second course of action (Ans.2) will provide the highest net income.